# Learning-based Multi-View Stereo: A Survey

### v1

@@ II-H1 Online MVS
II-H2 Offline MVS
Based on the depth estimation strategy as discussed in Sec. II-E, loss functions can be mainly categorized into regression and classification. For methods [15, 36] that predict depth with soft argmax [60], (smooth) $L_{1}$ loss is usually adopted as the loss function, which is stated as:
@@ II-H2 Offline MVS
III Unsupervised & Semi-supervised Methods with Depth Estimation
Figure 10: Ambiguous supervision from photometric consistency assumption [43]. (a) Complicated lighting condition caused by varying camera exposures. (b) Invisibility of specific area caused by occlusion.
In this section, we introduce unsupervised learning-based MVS methods [129, 130, 43, 44, 42] and semi-supervised method [46]. Supervised MVS methods mentioned in Sec. II depend extensively on the availability of accurate ground truth depth maps obtained through depth-sensing equipment. This requirement not only complicates the data collection process, making it labor-intensive and costly but also restricts these methods to limited datasets and primarily indoor settings. To make MVS practical in more general real-world scenarios, it is vital to consider alternative unsupervised learning-based methods that can provide competitive accuracy compared to the supervised ones without any ground truth depth. Existing unsupervised methods are built upon the assumption of photometric consistency (Sec. III-A), which indicates that corresponding pixels from different views of the same 3D point share similar features. These approaches are categorized into end-to-end ones (Sec. III-B) and multi-stage ones (Sec. III-C). SGT-MVSNet [46] is the only semi-supervised method so far, and we introduce it in Sec. III-D. We also analyze the primary challenges in unsupervised MVS in Sec. III-B.
III-A Photometric Consistency Assumption
@@
The computation of the structural similarity loss occurs between a synthesized image and the initial reference image, aiming to uphold contextual congruity. More precisely, the assessment of contextual similarity often involves the utilization of the Structural Similarity Index (SSIM) proposed by [134], a metric commonly employed for quantifying contextual consistency. This metric is defined as follows:
@@
where $\nabla_{u}$ and $\nabla_{v}$ refer to the gradient along x and y axis, $\tilde{d}=d/\bar{d}$ is the mean-normalized inverse depth, $M$ represents the set of valid pixels in the reference image.
III-B End-to-end Unsupervised Methods
End-to-end methods [131, 42, 43, 135, 136] are those methods that are trained from scratch with the same input information as supervised methods (Sec. II), while not using ground truth depth for supervision. Instead, they usually adopt photometric consistency, structural similarity and smoothness constraint as parts of their loss terms. Unsup_MVS [131] proposes to inherit the supervision signal of view synthesis, as discussed in Sec. III-A, and dynamically select $X$ “best” (lowest loss) values out of $Y$ loss maps.
However, the bottleneck in unsupervised MVS is to find such accurate photometric correspondences. As illustrated in Fig. 10, in the real scenarios, non-Lambertian surfaces, varying camera exposures, and occlusions will make the assumption of photometric consistency invalid. Hence, such an invalid assumption will lead to “ambiguous supervision”.
To alleviate the ambiguous supervision, JDACS [42] introduces semantic consistency in addition to photometric consistency. It proposes to extract semantic features with a pre-trained network and create semantic classification map by applying non-negative matrix factorization. Then pixel-wise cross-view segmentation consistency is calculated. The difference here is that the semantic map is warped, instead of the image. In this way, the semantic classification maps are supervised by cross-entropy loss between the warped map $S_{i}$ and reference map $S_{0}$, both encoded as pixel-wise one-hot vector. However, the proposed semantic cross-view consistency loss is unstable to converge during training and such high-level semantic information excludes details and boundaries in MVS.
RC-MVSNet [43] digs into providing reliable and occlusion-aware supervision by introducing neural rendering. It takes the advantage of both strong representation ability of Neural Radiance Field [28] and strong generalization ability of cost volume. First, the unsupervised CasMVSNet [32] backbone in RC-MVSNet yields initial depth map supervised by photometric consistency. In addition, it generates depth priors for the Gaussian-Uniform mixture sampling in the rendering-consistency (RC) network. Then after volumetric rendering of the reference view, the depth map is supervised by an extra reference view synthesis loss and a depth rendering consistency loss. In this way, occlusion-aware neural rendering with mixture sampling is utilized to alleviate the invisibility phenomenon and ambiguous supervision from different lighting condition is solved by reference view synthesis.
ElasticMVS [135] introduces an architecture of part-aware patchmatch to address limitations in photometric loss-based geometry information, which tends to have missing data and artifacts in certain areas, particularly textureless regions. The proposed framework incorporates an elastic part representation to encode geometric details for guiding piecewise-smooth depth map prediction. The initial step involves the formal definition of part-aware representation. Building upon this learned representation, enhancements are proposed for both depth hypotheses propagation and evaluation within the context of part-aware patchmatch. Finally, the network is optimized through the incorporation of a dense contrastive loss for self-supervised training, coupled with the integration of a spatial concentration loss [137] designed to promote the isotropic isolation of all pixel embeddings.
CL-MVSNet [136] proposes a framework that aims to enhance proximity among positive pairs by ensuring contrastive consistency between a regular branch and two contrastive branches. The regular branch is a CasMVSNet [32] structure that has been used in  [43, 135, 42] as well, while contrastive branches incorporate an image-level branch and a scene-level branch. These branches enforce contrastive consistency guided by the confidence mask estimated from the regular branch. Additionally, the $L0.5$ photometric consistency is proposed to enhance the accuracy of the reconstruction.
@@ III-B End-to-end Unsupervised Methods
Note that these end-to-end methods are all trained from scratch without any pre-processing which saves training time and reduces the complexity of application in real scenarios.
III-C Multi-stage Unsupervised Methods
@@
Multi-stage methods require either pre-training of a specific module or pre-processing on the training data. These methods are built upon the idea of pseudo-label generation.
Self-supervised CVP-MVSNet [132] proposes to generate pseudo ground truth of depth by training the CVP-MVSNet [75] backbone with photometric consistency loss (Sec. III-A). The pseudo ground truth depth maps are refined with high-resolution images, filtered with cross-view depth consistency check and then fused in a point cloud [15]. Then the mesh is reconstructed with the screened Poisson surface reconstruction algorithm [67] to render complete pseudo depth maps for training in the next iteration. The whole pipeline consists of several iterations to improve the performance. However, [132] cannot provide a satisfying reconstruction of the 3D objects since it heavily relies on the photometric consistency assumption. Additionally, the self-training stage of the whole pipeline takes few days to complete.
U-MVSNet [44] introduces a flow-depth consistency loss by pre-training an optical-flow estimation network, PWC-Net [138], in an unsupervised manner. The dense 2D optical flow correspondences are used to generate pseudo labels for cross-view flow-depth consistency, which alleviates the ambiguous supervision in the foreground. Then the uncertainty-aware self-training consistency further reduces the invalid supervision in the background utilizing the generated pseudo labels and uncertainty map.
Recently, KD-MVS [45] achieves superior performance by leveraging a typical teacher-student scheme of knowledge distillation [139, 140, 141]. It trains a teacher MVS model, e.g., MVSNet [15] or CasMVSNet [32], in a self-supervised manner, by enforcing both photometric and feature-metric consistency between the reconstructed and original images. Then the teacher model produces pseudo ground-truth labels of the training set by cross-view consistency check and probabilistic encoding. In addition to the reliable estimation by the teacher model, the pseudo labels also contain probabilistic knowledge with estimation uncertainty. By training MVS networks [15, 32] with the pseudo labels, the knowledge of the teacher model is distilled to the student models.
III-D Semi-supervised Methods
@@
SGT-MVSNet [46] proposes to use only a few sparse ground truth 3D points to estimate the depth map of the reference view. 3D point consistency loss is used to minimizes the difference between the 3D points back-projected from the corresponding pixels and the ground truth. To tackle the problem of inaccurate estimation on the edge and boundary, a coarse-to-fine reliable depth propagation module rectifies the erroneous predictions.
IV Learning-based MVS without Depth Estimation
@@
IV-A Voxel-based Methods
These methods [47, 48, 142, 143] estimate the scene geometry with volumetric representation by leveraging implicit function, e.g., SDF. Specifically, Atlas [47] and NeuralRecon [48] attempt to predict the TSDF volume from the 3D feature volume constructed by lifting 2D image features. As shown in Fig. 11, Atlas uses 3D CNN to regress the TSDF volume based on the feature volume accumulated from all images of the scene, which exhibits great completeness of reconstruction. The TSDF reconstruction is supervised using $L_{1}$ loss to the ground truth TSDF vaules. Since dense feature volume brings lots of computational overhead, NeuralRecon further improves the efficiency by incrementally reconstructing the scene in a fragment-wise and coarse-to-fine manner. The 3D features from different fragments are passed through the whole incremental reconstruction process with a RNN. TransformerFusion [142] fuses coarse and fine image features in a voxel grid with two transformers and then predicts an occupancy field to represent the scene geometry. VoRTX [143] uses a similar design to TransformerFusion and the scene geometry is obtained by passing three different-level features output from transformers through a 3D CNN.
Figure 11: Pipeline of Atlas [47]. 2D image features are back-projected into 3D volumes, which are aggregated and passed through a 3D CNN to directly regress a TSDF volume.
IV-B NeRF-based Methods
In novel view synthesis, Neural Radiance Field (NeRF) [28] has kicked off a new emerging representation of 3D, which offers a differentiable volume-rendering scheme to supervise a 3D radiance-based representation with 2D image-level losses. The pipeline of NeRF is shown in Fig. 12. NeRF employs multi-layer perceptron (MLP) to map a position $(x,y,z)$ and the normalized view direction $(\theta,\phi)$ to the corresponding color $\mathbf{c}$ and volume density $\sigma$. For a specific ray at a novel viewpoint, NeRF uses approximated numerical volume rendering to compute the accumulated color as:
@@ IV-B NeRF-based Methods
Figure 12: Pipeline of NeRF [28]. Given a 3D position and 2D viewing direction (a), an MLP produces the color and volume density (b). Then volume rendering is used to composite these values into an image (c). The optimization is minimizing the rendering loss (d).
@@
IV-C 3D Gaussian Splatting-based Methods
@@
Motivated by the NeRF-based MVS methods, researchers try to adapt 3DGS for reconstruction. Though 3DGS achieves high-quality novel-view synthesis, it is challenging to recover high quality geometry since no explicit geometry constraint is used and 3D Gaussians do not correspond well to the actual surface because of the 3D covariance [52]. As one of the earliest methods, DreamGaussian [51] follows NeRF [28] to compute a dense density grid with the 3D Gaussians and then extracts mesh with Marching Cubes [66]. SuGaR [52] introduces a geometry regularization term encouraging the 3D Gaussians to be well-aligned over the scene surfaces so that the Gaussians can contribute to better scene geometry. To reconstruct the mesh, SuGaR samples 3D points on a level set of the density computed from 3D Gaussians and then runs Poisson Reconstruction [172] on these points. NeuSG [173] jointly trains Neuralangelo [158], a NeRF-based method, and 3DGS. During training, the normals from the SDF field of Neuralangelo regularize the rotation of 3D Gaussians, while the SDF field is regularized to ensure that the SDF values at 3D Gaussians’ positions are close to zero. However, retrieving good geometry from Gaussian-based representations still needs to be explored.
IV-D Large Feed-forward Methods
In 3D reconstruction and 3D generation, a recent prominent trend is to directly learn the 3D representation from large-scale 3D datasets, such as Objaverse [30]. These methods typically adopt large-scale transformers [174] to generate a 3D representation based on various inputs, including single image [53, 175, 176], textual description [177], posed multi-view images [178], and un-posed multi-view images [179, 54]. LRM [53] uses a transformer-based feed-forward network to regress the 3D object from a single image. Specifically, LRM predicts features of a tri-plane representation [180], which is subsequently converted into an implicit field like NeRF [53]. Then the surface is extracted with Marching Cubes [66].
Recently, DUSt3R [54] reconstructs the scene as a point cloud given a pair of un-calibrated and un-posed images. After extracting image token representations with Vision Transformer [181], two transformer decoders exchange information of the token representations via cross-attention and then regress the per-pixel 3D points in the coordinate frame of the first image. Many downstream tasks can be performed with the point clouds, e.g., point matching, localization, intrinsic and relative pose estimation.
V Discussions
@@
In this section, we summarize and discuss the performance of learning-based MVS methods, including supervised online methods with depth estimation (Sec. V-A), supervised offline methods with depth estimation (Sec. V-B), unsupervised methods with depth estimation (Sec. V-C) and methods without depth estimation (Sec. V-D). Moreover, we discuss potential directions for future research (Sec. V-E).
@@
TABLE II: Quantitative results of supervised online MVS methods on ScanNet [26] and 7-Scenes [27].
@@
abs $\downarrow$ abs-rel $\downarrow$ abs-inv $\downarrow$ $\eta<1.25$uparrow abs $\downarrow$ abs-rel $\downarrow$ abs-inv $\downarrow$ $\eta<1.25$uparrow
MVDepthNet [25] 0.167 0.087 0.054 0.925 0.201 0.117 0.071 0.877
DPSNet [97] 0.219 0.119 0.071 0.868 0.249 0.149 0.085 0.826
Neural-RGBD [87] 0.236 0.122 0.075 0.850 0.214 0.131 0.076 0.865
GP-MVS [58] 0.149 0.076 0.049 0.940 0.174 0.100 0.064 0.903
Long et al. [94] 0.151 0.081 - 0.931 0.253 0.147 - 0.804
DeepVideoMVS [31] 0.119 0.060 0.038 0.965 0.145 0.038 0.054 0.938
MaGNet [88] 0.147 0.081 - 0.930 0.213 0.126 - 0.855
RIAV-MVS [111] 0.139 0.075 - 0.938 0.178 0.100 - 0.897
SimpleRecon [35] 0.089 0.043 - 0.981 0.105 0.058 - 0.974
TABLE III: Quantitative results of point cloud evaluation on MVS benchmarks [21, 22, 23] for unsupervised depth map-based MVS methods.“Finetuned with BlendedMVS” denotes whether the methods are finetuned on BlendedMVS [24] before evaluating on Tanks and Temples [22].
@@
TABLE IV: Quantitative results of point cloud evaluation on MVS benchmarks [21, 22, 23] for supervised offline MVS methods. “Finetuned with BlendedMVS” denotes whether the learning-based methods are finetuned on BlendedMVS [24] before evaluating on Tanks and Temples [22] and ETH3D [23].
@@
V-A Supervised Online MVS with Depth Estimation
Typically, online MVS methods are trained on ScanNet [26] and then evaluated on ScanNet [26] and 7-Scenes [27]. We summarize the quantitative results of online MVS methods in Tab. II. By incorporating temporal information in depth estimation,  [58, 94, 31] achieve better performance than MVDepthNet [25]. However, this usually increase the network complexity. Instead of temporal fusion, SimpleRecon [35] injects cheaply available metadata into the cost volume and achieves state-of-the-art performance on both ScanNet and 7-Scenes.
V-B Supervised Offline MVS with Depth Estimation

### v2

@@ 3.8.1 Online MVS
3.8.2 Offline MVS
Based on the depth estimation strategy as discussed in Sec. 3.5, loss functions can be mainly categorized into regression and classification. For methods [15, 35] that predict depth with soft argmax [62], (smooth) $L_{1}$ loss is usually adopted as the loss function, which is stated as:
@@ 3.8.2 Offline MVS
4 Unsupervised & Semi-supervised Methods with Depth Estimation
In this section, we introduce unsupervised learning-based MVS methods [125, 126, 42, 43, 41] and semi-supervised method [45]. Supervised MVS methods mentioned in Sec. 3 depend extensively on the availability of accurate ground truth depth maps obtained through depth-sensing equipment. To make MVS practical in more general real-world scenarios, it is vital to consider alternative unsupervised learning-based methods that can provide competitive accuracy compared to the supervised ones without any ground truth depth. Existing unsupervised methods are built upon the assumption of photometric consistency (Sec. 4.1), and are categorized into end-to-end (Sec. 4.2) and multi-stage (Sec. 4.3). SGT-MVSNet [45] is the only semi-supervised method so far, and we introduce it in Sec. 4.4.
4.1 Photometric Consistency Assumption
@@
The computation of the structural similarity loss occurs between a synthesized image and the initial reference image, aiming to uphold contextual congruity. More precisely, the assessment of contextual similarity often involves the Structural Similarity Index (SSIM) proposed [130], defined as:
@@
where $\nabla_{u}$ and $\nabla_{v}$ refer to the gradient along x and y axis, $\tilde{d}=d/\bar{d}$ is the mean-normalized inverse depth, and $M$ represents the set of valid pixels in the reference image.
4.2 End-to-end Unsupervised Methods
End-to-end methods [127, 41, 42, 131, 132] train from scratch with the same input as supervised methods (Sec. 3) but without ground truth depth. They are based on photometric consistency, structural similarity, and smoothness constraints for loss terms.
Unsup_MVS [127] adopts view synthesis supervision and dynamically selects the $X$ best (lowest loss) values out of $Y$ loss maps. However, photometric correspondences can be inaccurate due to non-Lambertian surfaces, camera exposure variations, and occlusions, leading to ambiguous supervision (Sec. 4.1). JDACS [41] introduces semantic consistency to address these challenges, using a pre-trained network to generate semantic maps and enforcing cross-view segmentation consistency. However, this approach can struggle to converge and lacks detailed information. RC-MVSNet [42] enhances supervision through neural rendering by combining Neural Radiance Fields and cost volumes. It uses a CasMVSNet backbone to generate depth maps supervised by photometric consistency, along with rendering-consistency and view synthesis losses to handle occlusions and varying lighting. ElasticMVS [131] addresses issues in photometric loss-based geometry by introducing a part-aware patch-match framework, using an elastic part representation to guide the depth map prediction. The network is optimized with contrastive and spatial concentration losses to promote pixel isolation. CL-MVSNet [132] improves proximity between positive pairs through contrastive consistency between a regular CasMVSNet branch and two contrastive branches. It also introduces $L0.5$ photometric consistency to improve the precision of the reconstruction.
@@ 4.2 End-to-end Unsupervised Methods
These end-to-end methods train from scratch without pre-processing, reducing training time and complexity in real-world applications.
4.3 Multi-stage Unsupervised Methods
@@
Multi-stage methods require pre-training or data pre-processing and are based on pseudo-label generation.
Self-supervised CVP-MVSNet [128] generates pseudo ground truth depth using photometric consistency, refining depth maps with cross-view consistency checks and point cloud fusion [15]. The meshes are then reconstructed for further training iterations. U-MVS [43] pre-trains an optical-flow network, PWC-Net [133], and uses flow-depth consistency to generate pseudo labels, reducing supervision ambiguity in foregrounds and backgrounds with an uncertainty-aware self-training consistency. KD-MVS [44] employs knowledge distillation to increase performance. A teacher MVS model generates pseudo ground-truth labels through photometric and feature-metric consistency loss, with uncertainty encoding. These labels are then used to train the student models.
4.4 Semi-supervised Methods
@@
SGT-MVSNet [45] uses sparse 3D points to estimate depth maps. A 3D point consistency loss minimizes differences between back-projected 3D points and ground truth. A coarse-to-fine depth propagation module improves accuracy at edges and boundaries.
5 Learning-based MVS without Depth Estimation
@@
5.1 Voxel-based Methods
These methods [46, 47, 134, 135] estimate the scene geometry with volumetric representation by leveraging implicit function, e.g., SDF. Specifically, Atlas [46] and NeuralRecon [47] attempt to predict the TSDF volume from the 3D feature volume constructed by lifting 2D image features. Atlas uses 3D CNN to regress the TSDF volume based on the feature volume accumulated from all images of the scene, which exhibits great completeness of reconstruction. The TSDF reconstruction is supervised using $L_{1}$ loss to the ground truth TSDF vaules. Since dense feature volume brings lots of computational overhead, NeuralRecon further improves the efficiency by incrementally reconstructing the scene in a fragment-wise and coarse-to-fine manner. The 3D features from different fragments are passed through the whole incremental reconstruction process with a RNN. TransformerFusion [134] fuses coarse and fine image features in a voxel grid with two transformers and then predicts an occupancy field to represent the scene geometry. VoRTX [135] uses a similar design to TransformerFusion and the scene geometry is obtained by passing three different-level features output from transformers through a 3D CNN.
5.2 NeRF-based Methods
In novel view synthesis, Neural Radiance Field (NeRF) [27] has kicked off a new emerging representation of 3D, which offers a differentiable volume-rendering scheme to supervise a 3D radiance-based representation with 2D image-level losses. NeRF employs multi-layer perceptron (MLP) to map a position $(x,y,z)$ and the normalized view direction $(\theta,\phi)$ to the corresponding color $\mathbf{c}$ and volume density $\sigma$. For a specific ray at a novel viewpoint, NeRF uses approximated numerical volume rendering to compute the accumulated color as:
@@ 5.2 NeRF-based Methods
5.3 3D Gaussian Splatting-based Methods
@@
Motivated by the NeRF-based MVS methods, researchers try to adapt 3DGS for reconstruction. Though 3DGS achieves high-quality novel-view synthesis, it is challenging to recover high quality geometry since no explicit geometry constraint is used and 3D Gaussians do not correspond well to the actual surface because of the 3D covariance [51]. SuGaR [51] introduces a geometry regularization term encouraging the 3D Gaussians to be well-aligned over the scene surfaces so that the Gaussians can contribute to better scene geometry. To reconstruct the mesh, SuGaR samples 3D points on a level set of the density computed from 3D Gaussians and then runs Poisson Reconstruction [162] on these points. NeuSG [163] jointly trains Neuralangelo [149], a NeRF-based method, and 3DGS. During training, the normals from the SDF field of Neuralangelo regularize the rotation of 3D Gaussians, while the SDF field is regularized to ensure that the SDF values at 3D Gaussians’ positions are close to zero. 2DGS [52] further improves surface accuracy by modeling the 3D primitives as surfels ( tangent plane clipped) and fuses rendered depth maps into mesh with TSDF fusion. PGSR [53] elicits unbiased depth maps as strong geometry prior from rendered distances and rendered normals, and enforces multi-view geometric consistency by regularizing such a prior with local plane assumption and edge constraint. RaDe-GS [164] also improves the quality of geometry by precisely modeling the depth and normal, instead of the error-prone approximate projection in the original 3DGS. However, retrieving good geometry from Gaussian-based representations still needs to be explored.
5.4 Large Feed-forward Methods
In 3D reconstruction and 3D generation, a recent prominent trend is to directly learn the 3D representation from large-scale 3D datasets, e.g.Objaverse [29]. These methods typically adopt large-scale transformers [165] to generate a 3D representation based on various inputs, including single image [54, 166, 167], textual description [168, 169], posed multi-view images [168, 170, 171, 172, 173], and un-posed multi-view images [174, 55]. LRM [54] uses a transformer-based feed-forward network to regress the 3D object from a single image. Specifically, it predicts features of a tri-plane representation [175], which is subsequently converted into a radiance field similar to NeRF. Large feed-forward models can also be combined with 3D Gaussian Splatting, where they directly generate 3D Gaussian points from posed multi-view inputs [168, 170, 171, 172, 173, 176]. Instead of predicting 3D Gaussian locations directly, which is a highly ill-posed task, the common practice is to use per-pixel aligned Plücker ray embeddings and predict the depth for each pixel [176] of input views. As such, this design requires the input views to cover a large range of the object. Recently, DUSt3R [55] predicts pixel-aligned point coordinates directly from uncalibrated image pairs. After extracting image token representations with Vision Transformer [177], two transformer decoders exchange information of the token representations via cross-attention and then regress the per-pixel 3D points in the coordinate frame of the first image. Many downstream tasks can be performed with the point clouds, e.g., point matching, localization, intrinsic and relative pose estimation. Directly extended from DUSt3R, MASt3R [56] focuses on the specific application of feature description and matching across multi-view images. It leverages an additional matching head on top of DUSt3R’s architecture to boost the performance of feature-based matching. With the dense correspondences established, MASt3R manages to achieve a comparable reconstruction quality as MVSNet [15]. Spann3R [178] manages an external spatial memory to keep track of all previous relevant 3D information yielded by DUSt3R and then queries this spatial memory to predict the 3D structure.
6 Discussions
@@
In this section, we summarize and discuss the performance of learning-based MVS methods, including supervised online methods with depth estimation (Sec. 6.1), supervised offline methods with depth estimation (Sec. 6.2), unsupervised methods with depth estimation (Sec. 6.3) and methods without depth estimation (Sec. 6.4). Moreover, we discuss potential directions for future research (Sec. 6.5).
@@
TABLE I: Quantitative results of supervised online MVS methods on ScanNet [26] and 7-Scenes [179].
@@
abs $\downarrow$ abs-rel $\downarrow$ $\eta<1.25$uparrow abs $\downarrow$ abs-rel $\downarrow$ $\eta<1.25$uparrow
MVDepthNet [25] 0.167 0.087 0.925 0.201 0.117 0.877
DPSNet [96] 0.219 0.119 0.868 0.249 0.149 0.826
Neural-RGBD [87] 0.236 0.122 0.850 0.214 0.131 0.865
GP-MVS [60] 0.149 0.076 0.940 0.174 0.100 0.903
DeepVideoMVS [30] 0.119 0.060 0.965 0.145 0.038 0.938
MaGNet [88] 0.147 0.081 0.930 0.213 0.126 0.855
RIAV-MVS [109] 0.139 0.075 0.938 0.178 0.100 0.897
SimpleRecon [34] 0.089 0.043 0.981 0.105 0.058 0.974
DoubleTake [89] 0.077 0.037 0.984 0.099 0.053 0.970
TABLE II: Quantitative results of point cloud evaluation on MVS benchmarks [21, 22, 23] for unsupervised depth map-based MVS methods.“Finetuned with BlendedMVS” denotes whether the methods are finetuned on BlendedMVS [24] before evaluating on Tanks and Temples [22].
@@
TABLE III: Quantitative results of point cloud evaluation on MVS benchmarks [21, 22, 23] for supervised offline MVS methods. “Finetuned with BlendedMVS” denotes whether methods are finetuned on BlendedMVS dataset [24] before evaluation on Tanks and Temples [22] and ETH3D [23].
@@
6.1 Supervised Online MVS with Depth Estimation
Typically, online MVS methods are trained on ScanNet [26] and then evaluated on ScanNet [26] and 7-Scenes [179]. We summarize the quantitative results of online MVS methods in Tab. I. By incorporating temporal information in depth estimation,  [60, 188, 30] achieve better performance than MVDepthNet [25]. However, this usually increase the network complexity. Instead of temporal fusion, SimpleRecon [34] injects cheaply available metadata into the cost volume and achieves state-of-the-art performance on both ScanNet and 7-Scenes.
6.2 Supervised Offline MVS with Depth Estimation